Let S be a finite set, with |S|= 200. Find the number of subsets of S containing more than 2 elements
1
Expert's answer
2020-11-16T19:52:32-0500

The number of subsets of "S" containing k elements is equal to "\left(\begin{array}{c}n\\k\end{array}\right)=\frac{n!}{k!(n-k)!}", where "n=|S|=200."


Firstly, find the number "m" of subsets of "S" containing not more than 2 elements. The emptyset is a unique set containg no elements. There are 200 different subsets of "S" having one element. The number of subsets containg two elements is "\left(\begin{array}{c}200\\2\end{array}\right)=\frac{200!}{2!\cdot198!}=\frac{200\cdot199\cdot198!}{2\cdot 198!}=100\cdot199=19900."

Then "m=1+200+19900=20101." Since the set "S" contains "2^{200}" subsets, the number of subsets of "S" containing more than 2 elements is "2^{200}-m=2^{200}-20101."